# Longteng Guo1,3, Jie Jiang1,3, Jing Liu1,3 1Zidongtaichu Foundation Model Research Center, CASIA 2State Key Laboratory of Multimodal Artificial Intelligence Systems, CASIA

### v1

@@ 3.3 Ablation Study
Table 4: Comparison of the models with the pre-training objective of MAE vs. AR on 5 downstream tasks. Default setting is highlighted in blue.
Method Loss ER MI MW SS CM
@@
Current EEG pre-training models employ a masked signal reconstruction framework using bidirectional attention (Jiang et al., 2024; Yang et al., 2024b). In this framework, random segments of the signal are masked, and the resulting input is processed by an encoder that reconstructs these segments based on contextual information (i.e., MAE). To enable a rigorous comparison between MAE and autoregressive (AR) modeling, we conduct an in-depth analysis using three distinct reconstruction loss functions: $\ell_{1}$, $\ell_{2}$, and cosine. For fairness, both pre-training paradigms utilize the same model architecture and parameter settings. We report related results in Table 4. For clarity, the standard deviations of the results presented have been omitted. The conclusions are twofold. First, for the three types of reconstruction loss, $\ell_{2}$ outperforms $\ell_{1}$, while $\cos$ shows the least efficacy. Second, irrespective of the distance metric employed, the AR architecture consistently outshines the MAE architecture, demonstrating a more than 2% average accuracy advantage. These findings align with and support existing research in the NLP domain (Radford et al., 2019; Achiam et al., 2023; Brown, 2020). Specifically, the unidirectional modeling task poses significant challenges, enabling the model to learn more robust representations. Additionally, AR effectively adapts to the temporal characteristics of EEG signals, capturing their patterns more directly and naturally.
Mutual enhancement is observed among various tasks.
Table 5: Comparison of the models with the settings of joint training vs. separate training on 5 downstream tasks. Default setting is highlighted in blue.
Settings ER MI MW SS CM
@@
We compare two downstream task training settings—joint multi-task training (default) and separate training—as shown in Table 5. For clarity, the standard deviations of the results presented have been omitted. In the separate training setting, each model is trained independently for each task, utilizing the same number of iterations and architectures as in the joint training scenario. Our observations indicate that models utilizing joint training consistently outperform those with separate training across all five tasks. Actually, for the two different datasets, the corresponding subgraphs share overlapping nodes. These shared nodes (i.e., electrodes) provide a form of data augmentation that benefits both datasets. This augmentation is particularly important for tasks with limited samples. For instance, task MW has a total size of only 4K, whereas task SS reaches 42K. The accuracy benefits of joint training are more pronounced for task MW compared to task SS (i.e., 3.9% for MW vs. 0.9% for SS). This enhancement suggests that despite originating from different tasks, signals from the same electrode exhibit shared patterns that can be effectively transferred. The introduction of the shared graph network effectively integrates and utilizes these shared patterns while also decoupling the differences between tasks. This phenomenon may provide an intriguing basis for future research on cross-task learning in EEG studies.
Generalized representational ability even on unseen data.
In this section, we explore a interesting conclusion regarding the transferability of ETE after autoregressive pre-training. Specifically, we utilize DREAMER (Katsigiannis & Ramzan, 2017), a dataset which is not included during the pre-training stage. This dataset comprises four categories formed by the $2\times 2$ combinations of high and low valence and arousal dimensions. It is fed into ETE to obtain representations. The entire process does not involve shared graph networks or require additional training. Consistent with the pre-training stage, for each electrode, we extract the last token as the global representation for that electrode. These last tokens are then averaged across electrode dimension, resulting in the final representation for each signal. We employ t-SNE (Van der Maaten & Hinton, 2008) to visualize the underlying structures and patterns within these representations, as illustrated in Figure 6. The findings indicate that autoregressive pre-training demonstrates strong transferability even on unseen data, effectively clustering signal from different patterns/categories together.
@@ 4 CONCLUSION
In conclusion, we have presented EEGPT, the first generalist EEG foundation model designed to overcome the limitations of existing specialized EEG models. By introducing an electrode-wise modeling strategy, developing an autoregressive pre-training approach, and implementing a multi-task transfer learning paradigm with a learnable electrode graph network, EEGPT unifies diverse EEG datasets and captures the sequential and temporal dependencies inherent in EEG signals. Our model demonstrates superior performance across 12 benchmarks, showcasing its versatility and scalability. We hope that EEGPT will inspire further research and development in generalist EEG models.
Appendix A Related Work
Despite the significant success of self-supervised pre-training in the fields of CV and NLP (Bai et al., 2024; Touvron et al., 2023), the potential of self-supervised pre-training in EEG remains underexplored. Specifically, existing work (Yang et al., 2024b; Jiang et al., 2024; Yi et al., 2024; Li et al., 2024) exhibits significant similarities in both the pre-training objective and the downstream task transfer paradigm. For pre-training objective, they predominantly employ the mask signal modeling (MAE) architecture. For instance, BIOT (Yang et al., 2024b) and MMM (Yi et al., 2024) adopt the channel and temporal embeddings to construct the EEG tokens for masked segments prediction. LaBraM (Jiang et al., 2024) utilizes a neural tokenizer to segment and encode EEG signals into discrete codes, and similarly predicts masked tokens from visible patches. However, the MAE architecture does not align with the temporal characteristics inherent in EEG signals. The models in the filed of NLP have shifted towards autoregressive models to better address similar temporal properties (Yang et al., 2024a; Touvron et al., 2023). Consequently, it is imperative to update the paradigm of self-supervised pre-training for EEG. Besides, these models encode EEG signals only for selected subsets of electrodes, which are lack of scalability and versatility. For downstream task transfer paradigm, they are still stuck to fine-tune separate models for each downstream task (Yang et al., 2024b; Jiang et al., 2024; Li et al., 2024) or even each subject (Yi et al., 2024), lacking the versatility required for an all-in-one model to multiple EEG tasks. Moreover, current pre-training models primarily focus on seizure epilepsy classification or emotion recognition tasks. In constrast, EEGPT aims for broader task coverage, enhancing both generalizability and adaptability.
Appendix B Pre-training Dataset Description
The detailed introduction of the datasets that we use for pre-training in our work and the data preprocessing procedure are as follows:
  • FACED (Chen et al., 2023): FACED is a large finer-grained affective computing EEG dataset based on the discrete model, consisting of 30-channel EEG data recorded at a sampling rate of 250 or 1,000 Hz from 123 participants.
  • SEED (Zheng & Lu, 2015): SEED is an emotion recognition dataset based on the discrete model, consisting of 62-channel EEG data recorded at a sampling rate of 1,000 Hz from 15 participants.
  • SEED-FRA (Liu et al., 2022): SEED-FRA is an emotion recognition dataset based on the discrete model, consisting of 62-channel EEG data recorded at a sampling rate of 1,000 Hz from 8 French participants.
  • SEED-GER (Liu et al., 2022): SEED-GER is an emotion recognition dataset based on the discrete model, consisting of 62-channel EEG data recorded at a sampling rate of 1,000 Hz from 8 German participants.
  • SEED-IV (Zheng et al., 2018): SEED-IV is an emotion recognition dataset based on the discrete model, consisting of 62-channel EEG data recorded at a sampling rate of 200 Hz from 15 participants.
  • SEED-V (Liu et al., 2021): SEED-V is an emotion recognition dataset based on the discrete model, consisting of 62-channel EEG data recorded at a sampling rate of 200 Hz from 16 participants.
  • THINGS-EEG-10Hz (Grootswagers et al., 2022): THINGS-EEG-10Hz is a visual event-related potential (ERP) dataset that consists of 63-channel EEG data recorded at a sampling rate of 1,000 Hz from 50 participants. It includes 1,854 object concepts of 22,448 images from the THINGS (Hebart et al., 2019) stimulus set.
  • THINGS-EEG-5Hz (Gifford et al., 2022): THINGS-EEG-5Hz is a visual event-related potential (ERP) dataset that consists of 122-channel EEG data recorded at a sampling rate of 1,000 Hz from 10 participants. It includes 1,854 object concepts of 16,740 images from the THINGS stimulus set.
  • IMG (Private): IMG is a visual event-related potential (ERP) dataset that consists of 122-channel EEG data recorded at a sampling rate of 1,000 Hz from 32 participants. It includes five semantic categories of 2,500 images of the visual perception task.
For data preprocessing, all the EEG signals are resampled to 256 Hz. The signals are then filtered between 0.1 and 100 Hz and segmented into samples of four seconds. Each sample is further segmented into 25 tokens with an overlap rate of 0.875, while each token has 256 sampling points. Besides, we apply the z-score normalization. No further preprocessing or artifact correction methods are applied.
Appendix C Multi-task Dataset Description
The detailed introduction of the datasets that we use for downstream tasks in our work are as follows:
  • DEAP (Koelstra et al., 2011): DEAP is an emotion recognition dataset based on the dimensional model, consisting of 32-channel EEG data recorded at a sampling rate of 128 Hz from 32 participants. It describes emotion from two dimensions: valence and arousal, each comprising two categories—high and low. We employ a four-class classification based on these dimensions for the emotion recognition task.
  • FACED (Chen et al., 2023): FACED is a large finer-grained affective computing EEG dataset based on the discrete model, consisting of 30-channel EEG data recorded at a sampling rate of 250 or 1,000 Hz from 123 participants. It contains data for nine emotion categories: amusement, inspiration, joy, tenderness; anger, fear, disgust, sadness, and neutral emotion. We employ a nine-class classification for the emotion recognition task.
  • SEED-IV (Zheng et al., 2018): SEED-IV is an emotion recognition dataset based on the discrete model, consisting of 62-channel EEG data recorded at a sampling rate of 200 Hz from 15 participants. It contains data for four emotions: happy, sad, neutral, and fear. We employ a four-class classification for the emotion recognition task.
  • SEED-V (Liu et al., 2021): SEED-V is an emotion recognition dataset based on the discrete model, consisting of 62-channel EEG data recorded at a sampling rate of 200 Hz from 16 participants. It contains data for five emotions: happy, sad, disgust, neutral, and fear. We employ a five-class classification for the emotion recognition task.
  • MIBCI (Cho et al., 2017): MIBCI is a motor imagery dataset, consisting of 64-channel EEG data recorded at a sampling rate of 512 Hz from 52 participants. We employ a binary classification based on the left and right hands motor imagery.
  • BCI Competition IV-1 (BCIC4-1) (Blankertz et al., 2007): BCIC4-1 is a motor imagery dataset which contains 59 channels of EEG data at a 100Hz sampling rate of 7 participants. We employ a binary classification based on the left or right hands and the both feet motor imagery.
  • EEGMat (Zyma et al., 2019): EEGMat is a mental workload dataset comprising 23-channel EEG data recorded at a sampling rate of 500 Hz from 36 participants. The dataset includes two categories of states: rest and doing tasks. We employ a binary classification based on these states for the mental workload detection task.
  • STEW (Lim et al., 2018): STEW is a mental workload dataset that includes 14-channel EEG data recorded at a sampling rate of 128 Hz from 45 participants. The dataset encompasses three levels of mental workload: low, medium, and high, allowing us to employ a three-class classification for the mental workload detection task.
  • EDF (Kemp et al., 2000): The EDF dataset comprises 2-channel EEG data recorded at a sampling rate of 100 Hz from 78 participants. It includes five sleep stages: wake, N1, N2, N3, and movement, enabling us to conduct a five-class classification for the sleep stage.
  • HMC (Alvarez-Estevez & Rijsman, 2021): The HMC dataset is a sleep dataset that comprises 4-channel EEG data recorded at a sampling rate of 256 Hz from 151 participants. It includes five sleep stages—wake, N1, N2, N3, and REM—facilitating a five-class classification for the sleep stage classification task.
  • IMG (Private): IMG is a visual event-related potential (ERP) dataset that consists of 122-channel EEG data recorded at a sampling rate of 1,000 Hz from 32 participants. It includes five semantic categories of 2,500 images for a five-class classification of the visual perception task.
  • SPE (Nguyen et al., 2017): SPE is a speech imagery dataset that consists of 64-channel EEG data recorded at a sampling rate of 256 Hz from 7 participants. It includes two types of words—long ("cooperate") and short ("in") for a binary classification of the cross-modality speech imagery task.
  • DREAMER (Katsigiannis & Ramzan, 2017): DREAMER is an emotion recognition dataset based on the dimensional model that consists of 14-channels EEG data recorded at a sampling rate of 128Hz from 23 participants. It describes emotion from two dimensions: valence and arousal, each comprising two categories—high and low. We employ the four-class data for analysis.
Appendix D Hyperparameter Settings
In this section, we detail the training protocols of EEGPT. The specific hyper-parameter configurations for the Stage I: Autoregressive pre-training and the Stage II: Multi-task fine-tuning are reported in Table 6. The training time is based on 8 NVIDIA A800-80G GPUs.
Table 6: Training hyperparameters for EEGPT of two training stage.
Stage Hyperparameter Base Large Huge Giant
Stage I Lr 1e-4
Time 3.2h 7.6h 11.2h 19.8h
Epoch 3.0
Precision BF16
Deepspeed Zero2 Zero2 Zero2 Zero3
LR Schedule cosine decay
Warmup Ratio 0.03
Batch Size per GPU 512
Gradient Checkpoint True
Stage II Lr 1e-4
Time 0.3h 0.7h 1.6h 2.9h
Epoch 10
Precision BF16
Deepspeed Zero2 Zero2 Zero2 Zero2
LR Schedule cosine decay
Warmup Ratio 0.1
Batch Size per GPU 64
Gradient Checkpoint True
Appendix E Baseline Model Description
The detailed descriptions of the six baseline models that we reproduce for comparison in this work are as follows:
  • EEGNet (Lawhern et al., 2018): EEGNet is a compact convolutional neural network designed for EEG-based brain-computer interfaces. It leverages depthwise and separable convolutions to facilitate efficient feature extraction and classification.
  • TSception (Ding et al., 2022): TSception is a multi-scale convolutional neural network designed for EEG emotion recognition, capable of learning discriminative representations across both time and channel dimensions. The model incorporates a dynamic temporal layer to effectively capture dynamic temporal and frequency representations, while an asymmetric spatial layer is employed to learn discriminative global and hemisphere representations.
  • Conformer (Song et al., 2023): Conformer is a compact convolutional Transformer model designed for EEG classification, capable of encapsulating both local and global features. It incorporates a convolutional module to effectively learn low-level local features while employing a self-attention mechanism to extract global correlations within the local temporal features.
  • LGGNet (Ding et al., 2023): LGGNet is a neurologically inspired graph neural network designed for EEG representation learning. It effectively models the intricate relationships both within and between the brain’s functional regions.
  • BIOT (Yang et al., 2024b): BIOT is a self-supervised biosignal learning model that tokenizes biosignals of various formats into "sentences". It segments each channel separately into tokens and flatten the tokens to form "sentences". Through its MAE architecture, it can be pre-trained with unlabelled data.
  • LaBraM (Jiang et al., 2024): LaBraM is a self-supervised EEG model that enables cross-dataset learning by segmenting the EEG signals into channel patches. It adopts the Vector-quantized neural spectrum prediction to train a neural tokenizer that encodes EEG patches into compact neural codes. Through its MAE architecture, it can be pre-trained with unlabelled data.
Appendix F Name of the Supporting Electrodes
The name of the supporting electrodes of our EEGPT are as listed in Table 7.
PO12 CCP2H FFC5H OI1 PO7 CPPZ
TP7 PO2 FC3 FTT7H PPO8 CCP4H
P11 FCC5H FFC4H FP1 CPP2H FFT7H
P1 I2 AFF6H FZ PO4 FCC2H
F8 FT9 CP2 AF3 FCZ POO11H
FPZ F3 P8 FC2 F1 CCP3H
CP6 PO1 C1 AFZ C3 CB1
FTT8H POO12H TP9 I1 FP2 POO10H
CPP1H CPP4H TTP8H AFF5H PO10 POO9H
POO3 CP5 PO3 FC6 FTT9H PPOZ
TPP5H POO4 CB2 FT7 CPZ CP1
PPO1 CP3 CCP5H O2 FCC1H CP4
FT8 T9 PO5 P2 P5 POZ
FC1 CPP3H C5 P9 P10 PO6
FFT8H CCP1H C2 POOZ T7 POO7
FFC3H F6 FCCZ TPP8H F7 P4
P3 AF8 PPO2 AF4 FFC2H FFC1H
P6 F2 C6 P12 TP10 CZ
IZ CCP6H TP8 PO11 OI2 FC5
TTP7H CPP5H F5 POO8 CPP6H OZ
PO9 AF7 PZ O1 FC4 PO8
F4 FCC3H T10 P7 FT10 FCC4H
FCC6H T8 PPO7 C4 FFC6H FTT10H
Table 7: Supporting EEG electrodes.

### v2

@@ IV-E3 Autoregression outperforms bidirectional masked pre-training
Table V: Comparison of the models with the pre-training objective of MAE vs. AR on 5 downstream tasks. Default setting is highlighted in blue.
Method Loss ER MI MW $S$ CM
@@
IV-E4 Mutual enhancement is observed among various tasks
Table VI: Comparison of the models with the settings of joint training vs. separate training on 5 downstream tasks. Default setting is highlighted in blue.
Settings ER MI MW $S$ CM
@@ IV-E4 Mutual enhancement is observed among various tasks
We compare two downstream task training settings—joint multi-task training (default) and separate training—as shown in Table VI. For clarity, the standard deviations of the results presented have been omitted. In the separate training setting, each model is trained independently for each task, utilizing the same number of iterations and architectures as in the joint training scenario. Our observations indicate that models utilizing joint training consistently outperform those with separate training across all five tasks. Actually, for the two different datasets, the corresponding subgraphs share overlapping nodes. These shared nodes (i.e., electrodes) provide a form of data augmentation that benefits both datasets. This augmentation is particularly important for tasks with limited samples. For instance, task MW has a total size of only 4K, whereas task SS reaches 42K. The accuracy benefits of joint training are more pronounced for task MW compared to task SS (i.e., 3.9% for MW vs. 0.9% for SS). This enhancement suggests that despite originating from different tasks, signals from the same electrode exhibit shared patterns that can be effectively transferred. The introduction of the shared graph network effectively integrates and utilizes these shared patterns while also decoupling the differences between tasks. This phenomenon may provide an intriguing basis for future research on cross-task learning in EEG studies.
IV-E5 Generalized representational ability even on unseen data
In this section, we explore an interesting conclusion regarding the transferability of ETE after autoregressive pre-training. Specifically, we utilize DREAMER [48], a dataset which is not included during the pre-training stage. This dataset comprises four categories formed by the $2\times 2$ combinations of high and low valence and arousal dimensions. It is fed into ETE to obtain representations. The entire process does not involve shared graph networks or require additional training. Consistent with the pre-training stage, for each electrode, we extract the last token as the global representation for that electrode. These last tokens are then averaged across electrode dimension, resulting in the final representation for each signal. We employ t-SNE [49] to visualize the underlying structures and patterns within these representations, as illustrated in Fig. 6. The findings indicate that autoregressive pre-training demonstrates strong transferability even on unseen data.
@@ V CONCLUSION
In conclusion, we have presented BrainGPT, the first generalist EEG foundation model designed to overcome the limitations of existing specialized EEG models. By introducing an electrode-wise modeling strategy, developing an autoregressive pre-training approach, and implementing a multi-task transfer learning paradigm with a learnable electrode graph network, BrainGPT unifies diverse EEG datasets and captures the sequential and temporal dependencies inherent in EEG signals. Our model demonstrates superior performance across 12 benchmarks, showcasing its versatility and scalability. We hope that BrainGPT will inspire further research and development in generalist EEG models.